# Training Machine Learning models at the Edge: A Survey

### v2

@@ V-E2 Comparison based on the requirements and needs of EL
Split Learning
Swarm Learning
Gossip Learning
@@
SNNs
Forward-Forward Algorithm
@@
\ : There is not enough information and literature to estimate the effect on the requirement
VI Edge Learning for different types of Machine Learning
@@
VI-A Unsupervised Learning
Considering the vast amount of unlabeled data produced in edge and end devices [32], it is very promising to use unlabeled data to train ML models on the edge. However, unsupervised learning comes with multiple challenges and restrictions for edge learning, especially when it comes to collaborative learning techniques such as FL or SL, which represent the vast majority of techniques used in the edge (see Figure 4). Unsupervised learning datasets may have a non-identically and independently distributed (non-IID) nature. Each node in a collaborative setting might have a different subset of the data, and the data distribution might vary across nodes. This non-IID property can make it difficult to effectively combine information from different nodes. Additionally, in the case of clustering algorithms, clusters may have varying sizes across nodes in a collaborative setting and clustering algorithms may need to adapt to changes in data distribution and cluster structures over time. Finally, since there are no available labels, and their assignment may differ between nodes (for example in a clustering algorithm). Ensuring consistency across distributed nodes is difficult, but crucial for aggregating meaningful global labels (clusters).
The figure 7 represents the main different types of unsupervised learning approaches used for training ML models in the edge, (a) using unsupervised learning algorithms directly on-device with non-collaborative methods [216, 35]; (b) using unsupervised learning methods to assist in the training of collaborative learning approaches, such as clustered federated learning [79, 217, 218] ; and (c) training an unsupervised learning model on the edge in a collaborative way [219].
(a) Unsupervised learning On a single device
(b) Unsupervised learning to assist collaborative learning
(c) collaborative unsupervised learning
Figure 7: Different types of Unsupervised Learning for Edge Learning
@@ VI-A Unsupervised Learning
Having an unsupervised learning trained on a single edge device is possible if the edge device has enough computation power and/or the learning algorithm is lightweight and can be trained with low resources, the training with such models is no different from the training on the cloud or other devices, the only difference being the constraint of low resources and data available [35]. Examples of unsupervised learning algorithms trained with this approach include [34] that investigates the application of K-Means on mainstream controllers, and [220] that presents the first dedicated CycleGAN accelerator for energy-constrained mobile applications, achieving a higher throughput-to-area ratio and higher energy efficiency than a GPU. In [216], an unsupervised segmentation was proposed that can be executed on edge devices without the need of annotated data. While [221], proposes TMNet, an approach to solve unsupervised video object segmentation problem at the edge. Finally, [222] propose an FPGA based architecture for a Self-Organization Neural Network capable of performing unsupervised learning on input features from a CNN by dynamically growing neurons and connections in order to perform class-incremental lifelong learning for object classification in the edge.
@@
Collaborative learning approaches like FL are promising solution for training ML models in the edge, However, FL the most popular technique on the edge, face multiple challenges such as non-independent and identically collaborative (non-IID) data and uneven computing power [110] and yields suboptimal results when the local clients’ data distributions diverge [79]. To solve these issues, the usage of clustering alongside FL have been proposed multiple times to cluster edge servers with similar environmental data distributions [223]. Among these approaches, Clustered Federated Learning (CFL), introduced by [79] is widely adopted, and it works by exploiting geometric properties of the FL loss surface and grouping the client population into clusters with jointly trainable data distributions. Approaches that build upon CFL have also been proposed, and they include [217, 110, 218, 224]. Other approaches that differ from CFL [79] include [68] that propose a hierarchical over-the-air FL approach which utilizes intermediary servers to form clusters near mobile users; and HPFL-CN [223], a communication-efficient hierarchical personalized federated edge learning framework, that achieves this by utilizing complex network feature clustering to group together edge servers that exhibit similar environmental data distributions. Subsequently, personalized models are trained for each cluster using a hierarchical architecture, resulting in enhanced efficiency. More specifically, HPFL-CN incorporates Privacy-preserving Feature Clustering (PFC) to derive privacy-preserving low-dimensional feature representations for each edge server. This is achieved by mapping the environmental data onto various complex network domains, thereby accurately clustering edge servers with similar characteristics. And Finally, [57] uses unsupervised learning methods on the edge to distinguish between classes across different users, when new classes with completely unseen data distributions emerge on devices in a federated learning setting for audio classification.
@@
as explained in previous sections, collaborative unsupervised learning methods are hard to train using unsupervised learning. Nevertheless, several methodologies have surfaced, one of which is FedUL (federation of unsupervised learning) [219]. In this approach, the unlabeled data undergo a transformation process to become surrogate labelled data for each client. Following this, a modified model is trained through supervised FL. Eventually, the desired model is obtained by recovering it from the modified model. An alternative methodology, known as FedUReID, has been proposed by Zhuang et al. [225] as a federated unsupervised person ReID system. The purpose of this system is to train person ReID models without the use of any labels, all the while ensuring the preservation of privacy.
VI-B Reinforcement learning
@@
Reinforcement learning (RL) has been successfully applied in the past on different problems in areas such as robotics, recommendation systems, video games and automatic vehicles [226, 227], making RL a promising and interesting direction for edge learning. However, The training of RL models in resources constrained environments is often limited by high compute and memory requirements from gradient computations [228], making the application of RL in the context of edge learning challenging. Despite these challenges, multiple RL approaches have been proposed for training ML models in the edge. Among them, Federated Reinforcement Learning (FRL) [229, 230] is a promising approach that allow multiple RL agents to learn optimal control policies for a series of devices with slightly different dynamics [231]. FRL is employed to achieve diverse objectives including personalization [232, 233], IoT traffic management [234, 235], Autonomous Systems [231] and resource allocation for unmanned aerial vehicle (UAV) [236]. FedMC [95] integrates RL models trained by multiple edge devices into a general model based on a meta-learning approach. FedGPO is an RL-based aggregation technique for FL introduced in [54], and aims to optimize the energy-efficiency of FL while guaranteeing model convergence. On the other hand, [237] introduces DDQN-Trust, a trust-based double deep Q-learning-based selection algorithm for FL that takes into account the trust scores and energy levels of the IoT devices to make appropriate scheduling decisions and integrate it with the main FL aggregation techniques (FedAvg, FedProx, FedShare and FedSGD). Other Federated Reinforcement Learning methods include [75, 238, 239, 240, 241, 242]
@@
Other Reinforcement learning approaches that don’t rely on FL or other collaborative learning paradigms have been proposed, such as [243] that uses online sequential learning to achieve full on-device reinforcement learning on an FPGA platform. In [244], a method using a combination of supervised learning and reinforcement learning was proposed for adaptive video streaming on edge servers or on-device, and [245] introduces an on-device RL-based adaptive video transmission algorithm to predict heterogeneous network bandwidth. EdgeRL, a lightweight C/C++ RL framework, was introduced in [246] and aims to provide a platform to train RL on edge devices. Finally, RL has also been employed in edge learning via shielding techniques [247], as proposed in [248] with a multi-agent system that enables each edge node to schedule its own jobs using SROLE, a shielded reinforcement learning technique used to check for action collisions that may occur because of the absence of coordination between the nodes, and provides alternative actions to avoid them.
VI-C Semi-supervised learning
As mentioned in the section VI-A using unlabeled data to train on the edge is very promising considering the vast amount of unlabeled data in edge and end devices [32]. Semi-supervised learning, constituting a machine learning paradigm, seeks to leverage both labeled and unlabeled data for specific learning tasks [249], presenting a means to harness the vast amount of unlabeled data at the edge, circumventing the challenges associated with supervised and unsupervised learning. Semi-supervised learning also provides other advantage on the edge, as training datasets are often not completely defined before training and might need to be completed using real-time data [250]. Naturally, several methodologies have surfaced to adapt semi-supervised learning on the edge, using federated learning based approaches [251, 252], or other learning techniques [253, 254, 255, 211, 256, 250, 250, 257].
VI-D Self-supervised learning
@@
Self-supervised learning (SSL) is a machine learning approach that allows models to learn from vast amounts of unlabeled data. SSL defines a pretext task based on unlabeled inputs, such as predicting the context surrounding a word in natural language or predicting masked patches of an image in computer vision, and learn generic representations that are useful across multiple tasks, and can be especially beneficial in domains where labelled data is scarce, or the specific task can not be known a priori [258].
@@
For this reason, Self-supervised learning is considered as a promising approach to leverage the abundance of unlabeled data at the edge for various AI applications. By learning from unlabeled data without explicit labels, self-supervised learning models can learn useful representations and skills that can be fine-tuned for specific tasks, such as recommendation systems [156, 148], speech and audio-related applications [259, 260], and others [261, 262, 263, 264]. Moreover, self-supervised learning can be particularly effective in scenarios where obtaining high-quality labels is challenging or expensive and in situation with data and concept drifting [261]. Several recent studies have proposed innovative self-supervised learning methods tailored for edge devices, such as contrastive learning [261, 263, 264].
VII Edge Learning Use Cases and applications
@@
As explained in previous sections, Edge Learning (EL) offer multiple advantages that range from low latency, bandwidth efficiency to privacy preservation and improved reliability and robustness, it also allows more customization and personalization by adapting to user preferences and behavior without relying on centralized computing or needing to collect and store private user data in cloud servers. In this section, we explore some uses cases and applications for EL that have been researched and developed in the past years.
VII-A Healthcare and Remote Monitoring
The use of ML in healthcare has been on constant improvement over the last years [265], despite this, cloud-based ML for healthcare continue to suffer from their limitations in meeting stringent security [266], privacy concerns [267], and low latency requirements [268]. As EL makes it possible to alleviate some of these challenges, it has gained considerable traction in the field of health care [45, 269, 270, 271]. Privacy concerns, being amongst the biggest challenges EL aims to solve for machine learning in healthcare [272, 45, 238]. Researches that aim to use EL for healthcare span in most of the field, from atrial fibrillation recognition [273], preterm labour risk prediction [274], cardiac disease detection [158], breast ultrasound image classification [255] to dermatological disease [264] and COVID-19 diagnosis, either using Clustered Federated Learning (CFL) [269] or blockchain technology and federated transfer learning on wearable devices [275].
VII-B Smart Technologies
@@
Smart cities are urban ecosystems designed using IoT technologies to solve urban life problems and improve the residents’ quality of life [279]. In this context, edge learning has been proposed to solve different challenges, mainly for its ability to leverage ML capabilities while preserving network bandwidth and reducing the charge on cloud servers. In [280], a cloud-aided edge learning based on knowledge fusion for smart lighting system has been proposed. Another application of ML in the edge is in smart grid systems where ML is needed to improve demand forecasting and automated demand response, as well as to analyse data related to energy use and obtain energy consumption patterns [281], detect anomalies [282], improve communications [283] and security [284] in the system. Other applications in smart cities include the detection of abnormal and dangerous activities [285, 286], pedestrian detection [287], water consumption forcasting [42, 288] and reducing congestions in intelligent traffic systems [289, 290]. Other more general EL approaches for smart cities include [291, 292, 293, 276, 252].
@@
Smart farming is another domain where ML is increasingly used to enhance the production quality, crop selection, and mineral deficiency detection, as well as to increase farmers’ earnings [294]. In [277] a TinyML based framework using deep neural networks and LSTM models for unmanned aerial vehicle (UAV) assisted smart farming was proposed, which measure soil moisture and ambient environmental conditions. Smart farming remains a promising domain for edge learning, although further research is needed for effectively harnessing its potential. Finally, using edge learning in smart homes can also be promising, however, at the time of writing this article, only a few papers explore this area, including [278].
VII-C Autonomous vehicles

### v3

@@ III-E2 Comparison based on the requirements and needs of edge learning
Split Learning \
Swarm Learning \ \ \ \
Gossip Learning \ \
@@
SNNs \ \ \
Forward-Forward Algorithm \ \
@@
\: There is not enough information and literature to estimate the effect on the requirement
Overall, analyzing the results from Table II, we can observe some high level trends, distributed techniques like FL and split learning exhibit computational efficiency due to their inherent distributed nature, reducing the load on individual devices. In contrast, we observe differences in memory footprint, with FL performing poorly due to its requirement for loading the entire model on each device, whereas split learning only requires loading a portion of the model. However, split learning is less efficient in terms of bandwidth usage and training time, as it necessitates frequent transmission of output from different splits. Similarly, gossip learning’s decentralized nature makes it less optimal for bandwidth usage, whereas swarm learning offers advantages in overcoming communication bottlenecks, reducing bandwidth usage. In terms of performance, FL often yields lower results than centralized alternatives, although this is not universally true across all approaches. It is worth noting that even when two techniques meet a requirement, they may not do so with the same level of efficiency. For instance, both FL and split learning meet the computational efficiency requirement, but split learning may offer greater efficiency, particularly when dealing with large models. Transfer learning, incremental learning, and meta-learning have all characteristics in common. Although they do not optimize for memory, as the model typically needs to be fully loaded for training, they often result in improved performance and reduced bandwidth usage compared to distributed methods. However, their impact on other requirements is generally less clear. Some approaches significantly optimize for these measures, while others don’t. Finally, model compression techniques inherently optimize for memory, computation, and energy requirements by reducing model complexity. As a result, they also reduce bandwidth usage compared to non-compressed models, simply by decreasing the model size. However, these methods often result in decreased performance.
To conclude with the analysis of Table II and Figures 3 and 4. Despite some drawbacks, FL has established itself as a cornerstone technique for edge learning, with successful adaptations across various domains and tasks. Moreover, combining FL with other techniques or specific implementations can mitigate its performance limitations and reduce memory and bandwidth usage. On the other hand, split learning shows great potential when combined with FL and is particularly promising for larger models, and we anticipate further advancements in this area. In contrast, adaptive and fine-tuning-based techniques are often a great choice when cloud pretraining is possible, reducing the amount of training needed on the edge, enabling further model personalization, and showing great potential when combined with distributed techniques. Model compression techniques are well-suited for edge devices, as they reduce model size, thereby decreasing computational, memory, and energy consumption. However, this often comes at the cost of decreased performance. Ultimately, each technique has its strengths, and the choice of technique should be based on the specific task and constraints at hand. Furthermore, combining multiple techniques can be beneficial, as it allows leveraging their individual strengths.
IV Edge Learning for different types of Machine Learning
@@
IV-A Unsupervised Learning
Considering the vast amount of unlabeled data produced in edge and end devices [45], it is very promising to use unlabeled data to train ML models on the edge. However, unsupervised learning comes with multiple challenges and restrictions for edge learning, especially when it comes to collaborative learning techniques such as FL or split learning, which represent the vast majority of techniques used in the edge (see Figure 3). Unsupervised learning datasets may have a non-IID nature. Each node in a collaborative setting might have a different subset of the data, and the data distribution might vary across nodes. This non-IID property can make it difficult to effectively combine information from different nodes. Additionally, in the case of clustering algorithms, clusters may have varying sizes across nodes in a collaborative setting and clustering algorithms may need to adapt to changes in data distribution and cluster structures over time. Finally, since there are no available labels, and their assignment may differ between nodes (for example in a clustering algorithm). Ensuring consistency across distributed nodes is difficult, but crucial for aggregating meaningful global labels (clusters).
Figures 56 and 7 represent the main different types of unsupervised learning approaches used for training ML models in the edge, (a) using unsupervised learning algorithms directly on-device with non-collaborative methods [228, 48] as shown in Figure 5; (b) using unsupervised learning methods to assist in the training of collaborative learning approaches, such as clustered federated learning [83, 229, 230] highlighted in Figure 6; and (c) training an unsupervised learning model on the edge collaboratively such as federated clustering [231] in Figure 7.
Figure 5: Unsupervised learning with the training happening only on a single device, with all the required data hosted locally
@@ IV-A Unsupervised Learning
Having an unsupervised learning model trained on a single edge device is possible if the device has enough computation power and/or the learning algorithm is lightweight and can be trained with low resources, the training with such models is usually no different from the training on the cloud or other devices, the only difference being the constraint of low resources and data available [48]. Examples of unsupervised learning algorithms trained with this approach include [47] which investigates the application of K-Means on mainstream controllers, and [232] that presents the first dedicated CycleGAN accelerator for energy-constrained mobile applications, achieving a higher throughput-to-area ratio and higher energy efficiency than a GPU. In [228], an unsupervised segmentation was proposed that can be executed on edge devices without the need of annotated data. While [233], proposes TMNet, an approach to solve unsupervised video object segmentation problem at the edge. Finally, [234] proposes an FPGA based architecture for a self-organization neural network capable of performing unsupervised learning on input features from a CNN by dynamically growing neurons and connections in order to perform class-incremental lifelong learning for object classification in the edge.
Figure 6: Unsupervised learning to assist collaborative learning, a clustering is typically applied to edge devices to improve on the FL process (For example CFL)
@@
Collaborative learning approaches, such as FL, are promising solutions for training ML models on edge devices. However, FL, the most popular technique on edge devices, faces multiple challenges, including non-IID data, uneven computing power [118] and suboptimal results when the local clients’ data distributions diverge [83]. To address these issues, the usage of clustering alongside FL have been proposed multiple times to cluster devices with similar environmental data distributions [235]. One popular approach is Clustered Federated Learning (CFL) [83], which exploits geometric properties of the FL loss surface and group the client into clusters with jointly trainable data distributions. CFL has been widely adopted and has inspired other approaches, such as [229, 118, 230, 236]. Alternative approaches include hierarchical over-the-air FL [237], which utilizes intermediary servers to form clusters near mobile users, and HPFL-CN [235], a communication-efficient hierarchical personalized FL framework that uses complex network feature clustering to group edge servers with similar environmental data distributions. Subsequently, personalized models are trained for each cluster using a hierarchical architecture, resulting in enhanced efficiency. HPFL-CN incorporates privacy-preserving feature clustering to derive low-dimensional feature representations for each edge server. This is achieved by mapping the environmental data onto various complex network domains, thereby accurately clustering edge servers with similar characteristics. Finally, [70] uses unsupervised learning methods on the edge to distinguish between classes across different users, when new classes with completely unseen data distributions emerge on devices in an FL setting for audio classification.
Figure 7: Collaborative unsupervised learning, the goal being to apply unsupervised learning algorithms on completely distributed data
@@
Collaborative unsupervised learning methods are challenging to train for all the reason explained previously. Nevertheless, several methodologies have surfaced. Among them is federated clustering, which aim to execute clustering on distributed data without sharing the data [231, 238, 239]. Federated clustering methods can be described as one-shot if they require only one round of transfer between clients and the servers [231, 240], or they can require multiple round of communication [241]. An alternative methodology, known as FedUReID, has been proposed by Zhuang et al. [242] as a person Re-identification system without the use of any labels, all the while ensuring the preservation of privacy. Finally, federation of unsupervised learning [243] proposes a method where unlabeled data undergo a transformation process to become surrogate labelled data for each client. Following this, a modified model is trained through supervised FL. Eventually, the desired model is obtained by recovering it from the modified model.
IV-B Reinforcement learning
@@
Reinforcement learning (RL) has been successfully applied in the past on different problems in areas such as robotics, recommendation systems, video games and automatic vehicles [244, 245], making RL a promising and interesting direction for edge learning. However, The training of RL models in resources constrained environments is often limited by high compute and memory requirements from gradient computations [246], making the application of RL in the context of edge learning challenging. Despite these challenges, multiple RL approaches have been proposed for training ML models in the edge. Among them, federated RL [247, 248] is a promising approach that allow multiple RL agents to learn optimal control policies for a series of devices with slightly different dynamics [249]. Furthermore, it is employed to achieve diverse objectives including personalization [250, 251], IoT traffic management [252, 253], Autonomous Systems [249] and resource allocation for unmanned aerial vehicle (UAV) [254]. FedMC [103] integrates RL models trained by multiple edge devices into a general model based on a meta-learning approach. FedGPO is an RL-based aggregation technique for FL introduced in [67], and aims to optimize the energy-efficiency of FL while guaranteeing model convergence. On the other hand, [255] introduces DDQN-Trust, a trust-based double deep Q-learning-based selection algorithm for FL that takes into account the trust scores and energy levels of the IoT devices to make appropriate scheduling decisions and integrate it with the main FL aggregation techniques (FedAvg, FedProx, FedShare and FedSGD). Other federated RL methods include [79, 256, 257, 258, 259, 260].
@@
Other RL approaches that don’t rely on FL or other collaborative learning paradigms have been proposed, such as [261] which uses online sequential learning to achieve full on-device RL on an FPGA platform. In [262], a method combining supervised and reinforcement learning is proposed for adaptive video streaming on edge servers or on-device, and [263] introduces an on-device RL-based adaptive video transmission algorithm to predict heterogeneous network bandwidth. Finally, RL has also been employed in edge learning via shielding techniques [264], as proposed in [265] with a multiagent system that enables each edge node to schedule its own jobs using SROLE, a shielded RL technique used to check for action collisions that may occur because of the absence of coordination between the nodes, and provides alternative actions to avoid them.
IV-C Semi-supervised learning
Semi-supervised learning, is a paradigm that combines labeled and unlabeled data for specific learning tasks [266]. It can be described as a middle ground between supervised and unsupervised learning, and leverages the advantages of both approaches. By combining a small amount of labeled with a large amount of unlabeled data. This is particularly beneficial in an edge learning context, where vast amounts of unlabeled data are generated continuously by end devices [45], and where sending the data for labeling in the cloud is not always possible for privacy reasons. For a comprehensive overview of semi-supervised learning, refer to the following survey [266]. Semi-supervised learning presents an alternative mean to harness the vast amount of unlabeled data at the edge. Additionally, it offers other advantage on the edge, as training datasets are often incomplete before training and might need supplementation with real-time data [267]. Various methodologies have been developed to adapt semi-supervised learning to edge environments, including FL based approaches [268, 269], as well as other learning techniques [270, 271, 272, 223, 273, 267, 267, 274].
IV-D Self-supervised learning
@@
Self-supervised learning is an ML approach that allows models to learn from vast amounts of data without explicit labels [275]. It creates labels from the data itself by defining pretext tasks where the data provides its own supervision. For instance, in natural language processing, a common pretext task involve predicting the context surrounding a word or predicting a given word in a sentence given the previous word, also known as language modeling. In computer vision, a pretext task might involve predicting masked patches of an image. Self-supervised learning can be especially beneficial in domains where labelled data is scarce, or the specific task can not be known a priori [275]. For a deeper exploration on self-supervised learning we refer the reader to [275].
@@
Because of its independence from labeled data, self-supervised learning is considered as a promising approach for edge learning. By learning from data without explicit labels, it enables learning useful representations and skills that can be fine-tuned for specific tasks, such as recommendation systems [168, 160], speech and audio-related applications [276, 277], and others [278, 279, 280, 281]. Moreover, self-supervised learning can be particularly effective in scenarios with data and concept drifting [278]. Several recent studies have proposed innovative self-supervised learning methods tailored for edge devices, such as contrastive learning [278, 280, 281].
V Edge Learning Use Cases and applications
@@
As explained in previous sections, edge learning offer multiple advantages that range from low latency, bandwidth efficiency to privacy preservation and improved reliability and robustness, it also allows more customization and personalization by adapting to user preferences and behavior without relying on centralized computing or needing to collect and store private user data in cloud servers. In this section, we explore some uses cases and applications for edge learning that have been researched and developed in the past years.
V-A Healthcare and Remote Monitoring
The use of ML in healthcare has been under constant improvement over the last years [282]. However, cloud-based ML solutions still struggle to meet the sector’s stringent security requirements [283], address privacy concerns [284], and satisfy low latency requirements [285]. Edge learning has emerged as a promising solution to address these challenges, gaining traction in the field, mostly by using federated edge learning [51, 286, 287, 288]. However, while FL in healthcare is increasingly explored as a privacy-preserving approach, the training of these models is often done with large resource requirements [289, 290], making it hard to implement on edge devices. Moreover, for various tasks, medical data is collected and managed directly by large organizations such as hospitals and medical facilities [291]. For such tasks FL with more powerful clients with access GPUs may be preferred, since the privacy constraints in these scenarios often involves data not being shared outside the organization rather than the local device. Nevertheless, for tasks where the data shouldn’t leave medical edge devices and wearables, such as sensors and ECG devices, edge learning remains a viable and promising direction [292, 293]. Researches that use edge learning for healthcare span in most of the field, from atrial fibrillation recognition [294], preterm labor risk prediction [295], cardiac disease detection [170], breast ultrasound image classification [272] to dermatological disease [281] and COVID-19 diagnosis, leveraging techniques such as CFL [286] and federated transfer learning [296]. For a more comprehensive overview of FL and edge learning in healthcare, refer to the survey [289].
When it comes to monitoring for medical purposes, Human Activity Recognition (HAR) represents one of the most popular use cases. HAR refers to the automation of the identification and categorization of the various activities performed by humans and their interactions with the environment [297]. As personalization for HAR has been shown to improve the results and performance of these systems [298], training the model on the edge using incremental learning or meta learning approaches can help achieve that while increasing the privacy and reducing the bandwidth consumption. PILOTE [136] proposes an incremental learning-based approach for HAR, designed for edge devices with extremely limited resources and demonstrates reliable performance in mitigating catastrophic forgetting. In addition, [299] proposes a personalizable lightweight CNN model for HAR, as well as a training algorithm to find personalization-friendly parameters. With the objective of improving the accuracy after the personalization when dealing with a wide range of target users. ClusterFL [300], proposes a clustering-based FL approach for edge-based HAR. Finally, [301] presents an on-device deep learning approach for STM32 microcontrollers, which fine-tunes a CNN model for enhanced HAR personalization.
V-B Smart Technologies
@@
Smart cities are urban ecosystems designed using IoT technologies to solve urban life problems and improve the residents’ quality of life [305]. In this context, edge learning has been proposed to solve different challenges, mainly for its ability to leverage ML capabilities while preserving network bandwidth and reducing the charge on cloud servers. In [306], a cloud-aided edge learning based on knowledge fusion for smart lighting system has been proposed. Another application of ML in the edge is in smart grid systems where ML is needed to improve demand forecasting and automated demand response, as well as to analyze data related to energy use and obtain energy consumption patterns [307], detect anomalies [308], improve communications [309] and security [310] in the system. Other applications in smart cities include the detection of abnormal and dangerous activities [311, 312], pedestrian detection [313], water consumption forecasting [58, 314] and reducing congestions in intelligent traffic systems [315, 316]. Other more general edge learning approaches for smart cities include [317, 318, 319, 302, 269].
@@
Smart farming is another domain where ML is increasingly used to enhance the production quality, crop selection, and mineral deficiency detection, as well as to increase farmers’ earnings [320]. In [303] a TinyML based framework using deep neural networks and LSTM models for unmanned aerial vehicles assisted smart farming was proposed, which measure soil moisture and ambient environmental conditions. Smart farming remains a promising domain for edge learning, although further research is needed for effectively harnessing its potential. Finally, using edge learning in smart homes can also be promising, however, at the time of writing this article, only a few papers explore this area, including [304].
V-C Autonomous vehicles